What is the pH of a solution made by mixing 30.0 mL solution of 0.313 M NaOH with 30.0 mL of 0.223 M HCl?
Answer #1

Molarity of mixer = ((30*0.313)-(30*0.223))/(30+30) = 0.045 M

as HCl is excess.

pH = -log(H+) = -log(0.045) = 1.35